OpenStudy (anonymous):

Find the area under the curve y=sqrt(4-x^2) that lies between x=1and x=2. Can someone tell me the steps needed to solve problems like this

zepdrix (zepdrix):

Ahhhh I'm so tired >.< I dunno if I have the energy for another math problemmmmm. Ok ok ok, let's try

OpenStudy (anonymous):

lol ok thank you

zepdrix (zepdrix):

It really helps if you're able to graph this, if not, it'll still work out ok. Do you understand what shape we get from that square root function?

OpenStudy (anonymous):

is the shape going to be half a circle?

zepdrix (zepdrix):

Top half of a circle? Ok good, y a :)

OpenStudy (anonymous):

can i solve it without graphing though...

zepdrix (zepdrix):

Ah, yes fine. Be that way. :3

zepdrix (zepdrix):

Since our function lies above the x-axis, it's simply:\[\Large\rm \int\limits \sqrt{4-x^2}dx\]And ya, we can throw the boundaries in there :)

zepdrix (zepdrix):

Do you need help dealing with this integral? Feels like a nice spot for a Trig Sub.

OpenStudy (anonymous):

would the boundaries be 1 and 2?

zepdrix (zepdrix):

Mmmm ya that seems about right,\[\Large\rm \int\limits\limits_1^2 \sqrt{4-x^2}dx\]

zepdrix (zepdrix):

This is kind of what it would look like ya?